# SIC-POVMs and orders of real quadratic fields

### v1

@@ Abstract.
We consider the problem of counting and classifying symmetric informationally complete positive operator-valued measures (SICs or SIC-POVMs), that is, sets of $d^{2}$ equiangular lines in $\mathbb{C}^{d}$. For $4\leq d\leq 90$, we show the number of known equivalence classes of Weyl–Heisenberg covariant SICs in dimension $d$ equals the cardinality of the ideal class monoid of (not necessarily invertible) ideal classes in the real quadratic order of discriminant $(d+1)(d-3)$. Equivalently, this is the number of $\mathbf{GL}_{2}(\mathbb{Z})$ conjugacy classes in $\mathbf{SL}_{2}(\mathbb{Z})$ of trace $d-1$. We conjecture the equality extends to all $d\geq 4$. We prove that this conjecture implies more that one equivalence class of Weyl–Heisenberg covariant SICs for every $d>22$.
Additionally, we refine the “class field hypothesis” of Appleby, Flammia, McConnell, and Yard [10] to predict the exact class field generated by the ratios of vector entries for the equiangular lines defining a Weyl–Heisenberg covariant SIC. The class fields conjecturally associated to SICs in dimension $d$ come with a natural partial order under inclusion; we show that the natural inclusions of these fields are strict, except in a small family of cases.
@@
symmetric informationally complete positive operator-valued measure (SIC-POVM), complex equianglular lines, equiangular tight frame (ETF), real quadratic field, non-maximal order, binary quadratic form, ideal class group, explicit class field theory, Stark conjectures
@@ 1. Introduction
This paper concerns provable and conjectural properties of SIC-POVMs, which are ensembles of quantum measurements described by $d^{2}$ equiangular complex lines in $\mathbb{C}^{d}$. These properties are number-theoretic and relate to class numbers and class fields attached to orders of associated real quadratic number fields.
@@
A set $S$ of complex lines through the origin in $\mathbb{C}^{d}$, written $S=\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{n}\}\subset\mathbb{P}^{d-1}(\mathbb{C})$, with $\mathbf{v}_{j}$ unit vectors, is equiangular if all the pairwise Hermitian inner products $\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle$ for $i\neq j$ have the same absolute value $|\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle|=\alpha\geq 0$. In 1975, Delsarte, Goethals, and Seidel [19] proved an upper bound of $n\leq d^{2}$ on the size of any set of complex equiangular lines in $\mathbb{C}^{d}$. They showed that if the bound $n=d^{2}$ is attained, then necessarily $\alpha=\frac{1}{d+1}$.
@@
A symmetric informationally complete positive operator-valued measure or SIC-POVM (henceforth SIC) is a set of quantum measurements that is equivalent to a (maximal) set of $d^{2}$ complex equiangular lines in $\mathbb{C}^{d}$. Specifically, for a set of $d^{2}$ equiangular lines $\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{d^{2}}\}$ represented by unit column vectors $\mathbf{v}_{j}\in\mathbb{C}^{d}$, the associated SIC is the set of normalized rank one Hermitian projection operators ${\mathbf{\Pi}}[S]:=\left\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\right\}$ where $\Pi_{j}=\mathbf{v}_{j}\mathbf{v}_{j}^{\dagger}$. (Vectors $\mathbf{v}$ are column vectors, $\mathbf{v}^{\dagger}$ is the conjugate transpose of $\mathbf{v}$, and the Hermitian inner product is $\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle=\mathbf{v}_{i}^{\dagger}\mathbf{v}_{j}$, following conventions from physics. Here, $\mathbf{v}_{i}^{\dagger}\mathbf{v}_{i}=1.$) The SIC condition is equivalent to the following $d^{4}$ identities on the traces of products of the projection operators:
@@
We will use the term “SIC” interchangeably to mean either a “set of $d^{2}$ (scaled) Hermitian projection matrices ${\mathbf{\Pi}}(S)$ forming a SIC-POVM” or a “set of $d^{2}$ complex equiangular lines $S$ in $\mathbb{C}^{d}$.” When more precision is needed, we term the latter a line-SIC.
The name SIC-POVM was introduced in a 2004 paper of Renes, Blume-Kohout, Scott, and Caves [36], whose study was motivated by quantum state tomography, the problem of reconstructing a quantum state from a series of measurements performed on identical states. Scott showed that SICs are optimal measurement ensembles for tomography [37]; see also [26]. SICs have found further applications to quantum information processing [14, 42, 45, 18], quantum foundations [25], compressed sensing for radar [28], and classical phase retrieval [21].
The object of this paper is to extend the range of conjectures about the structure of SICs proposed by Appleby, Flammia, McConnell, and Yard [9]. Their conjectures relate Weyl–Heisenberg SICs in $\mathbb{C}^{d}$ (defined in Section 2) for $d\geq 4$ to class field theory over real quadratic fields. The ratios of entries of each of the $d^{2}$ vectors in such SICs lie in a number field that is an abelian extension of $K=\mathbb{Q}(\sqrt{\Delta_{d}})$ with $\Delta_{d}=(d+1)(d-3)$. Class field theory classifies the abelian extensions of a given number field $K$ in terms of ray class groups, which encode the arithmetic of ideals of the ring $\mathcal{O}_{K}$ of algebraic integers of $K$.
The number fields associated to individual SICs in dimension $d$ may vary with the SIC, and the subclass of SICs to which the most precise conjecture of [9] applies consists of those giving rise to minimal number field that occurs for a fixed $d$. Appleby, Flammia, McConnell, and Yard conjecture that this minimal field is a specific ray class field of $\mathbb{Q}(\sqrt{\Delta_{d}})$; see 1.1. The corresponding SICs comprise the minimal multiplet, as defined in Section 2. It is one of several multiplets, which are finite sets of geometric equivalence classes of SICs making up a single orbit of a Galois group action on geometric equivalence classes.
The current paper presents a refinement of these conjectures that accounts for all the remaining multiplets of Weyl–Heisenberg SICs, those having larger associated number fields. For $d\geq 4$ the refined conjecture says that each multiplet is indexed by a particular order $\mathcal{O}^{\prime}$ of a real quadratic number field $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$ intermediate between $\mathcal{O}=\mathbb{Z}[\varepsilon_{d}]$ with $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$ and the maximal order $\mathcal{O}_{K_{d}}$. Here an order of a $K_{d}$ is a subring of $K_{d}$ whose additive group is free of rank $[K_{d}:\mathbb{Q}]=2$.
The number field associated to the $\mathcal{O}^{\prime}$-multiplet is conjectured to be a certain ray class field associated to the order $\mathcal{O}^{\prime}$. The definition of ray class fields associated to orders originates in the authors’ previous work [33]. Inclusions of orders induce reverse inclusion of the corresponding ray class fields, so the minimal multiplet corresponds to the maximal order.
1.1. SICs
@@
Most work on SICs follows Zauner’s 1999 PhD thesis [48, 49] on quantum designs. A few sporadic examples (such as the Hoggar lines in dimension $8$ [29, 30]) were known prior to Zauner’s work. Zauner conjectured, based on his numerical investigations, that SICs exist in every dimension $d$, and moreover, that a SIC in $\mathbb{C}^{d}$ could be found as the orbit of a single vector under the Weyl–Heisenberg group [48, 49]; the latter objects are called Weyl–Heisenberg covariant SICs (or just Weyl–Heisenberg SICs) and are defined and discussed in Section 2. Weyl–Heisenberg SICs are the main topic of this paper. (An extensive treatment of the history of SICs is given by Fuchs, Hoang, and Stacy [24].)
@@
Over the past 25 years, a large set of data has been computed about SICs in particular dimensions—sometimes approximately and sometime exactly—resulting in extensive datasets [39, 38, 22]. SICs have been rigorously constructed in a finite set of dimensions, including Weyl–Heisenberg SICs in all dimensions $1\leq d\leq 53$ and the Hoggar lines for $d=8$ (which are not Weyl–Heisenberg covariant). Beyond these exact constructions, numerical computations up to $d=90$ suggest there are only finitely many (equivalence classes of) Weyl–Heisenberg SICs in each dimension $d\neq 3$. It is known that $d=3$ has infinitely many equivalence classes of Weyl–Heisenberg SICs, and we will not consider it further.
@@
The current numerical data is believed to account for all Weyl–Heisenberg SICs of dimension $4\leq d\leq 90$. There is a geometric action moving these SICs around by the normalizer of the Weyl–Heisenberg group inside the extended unitary group (which is a finite group whose order varies erratically with $d$, while the Weyl–Heisenberg group itself has order $d^{3}$.) In addition, since the (normalized) generating vectors of Weyl–Heisenberg SICs are (empirically) given by algebraic numbers, there is a Galois action, which sometimes moves SICs to new SICs and sometimes moves them to non-SICs.
@@
    the finiteness of the set of (equivalence classes of) Weyl–Heisenberg SICs in every fixed dimension $d\neq 3$.
The finiteness assertion has been a “folklore” conjecture since about 2004 [2] and was observed informally up to dimension $d=50$ by Scott and Grassl (who state they are confident their list is complete for $d\leq 50$ in [39, Sec. 4]) and up to $d=90$ by Fuchs, Hoang, and Stacy [24, Sec. 6]. It was explicitly formulated as a conjecture in the first author’s PhD thesis [31, Conj. V.12]. A closely related conjecture was formulated by Waldron [46, Sec. 14.27].
The finiteness assertion would imply fiducial vectors of Weyl–Heisenberg SICs can be chosen to have all entries algebraic numbers. Further study of the algebraic numbers in existing examples has led to surprising discoveries.
@@ 1.2. Connection of SICs to algebraic number theory and class fields
Recent work has connected SICs to explicit class field theory over real quadratic number fields. Appleby, Yadsan-Appleby, and Zauner [4] discovered in 2013 that in all known examples, the algebraic number field $\mathbb{E}$ generated by the entries of the projections ${\mathbf{\Pi}}[S]$ contained the quadratic field $K_{d}=\mathbb{Q}(\sqrt{(d+1)(d-3)})$, and Galois automorphisms often mapped SICs to new SICs. Moreover, they observed that $\mathbb{E}$ was an abelian extension of $K_{d}$ in all known examples.
Such abelian Galois extensions are subfields of ray class fields over $K_{d}$. In 2016, Appleby, Flammia, McConnell, and Yard [10, 9] made explicit predictions relating the number fields generated by Weyl–Heisenberg SICs to ray class fields. Based on extensive data given in [10, Prop. 1], they formulated the following conjecture [10, Conj. 2].
@@
    Every Weyl–Heisenberg line-SIC $S$ in dimension $d$ has SIC-field $F^{\rm vec}\!\left(S\right)$ that is a finite Galois extension of $\mathbb{Q}$ containing $H_{d^{\prime}\infty_{1}\infty_{2}}$.
The technical definition of the SIC-field $F^{\rm vec}\!\left(S\right)$ appearing in 1.1 is given in Section 3.1. Ray class fields are discussed in Section 8.1.
@@
They verified assertion (2) of the conjecture for all currently known Weyl–Heisenberg SICs in these dimensions. They also verified that the fields $F^{\rm vec}\!\left(S\right)$ are all abelian over $K_{d}$ in the known cases. (The list of SICs used is believed to be complete in these dimensions.)
@@
The origin of the discriminant $\Delta_{d}=(d+1)(d-3)$ appearing in the conjecture is not completely understood, but there are some tantalizing hints. The appearance of the factor $d+1$ dividing $\Delta_{d}$ seems related to the fact that in the inner products of distinct unit vectors in a SIC are of absolute value $\frac{1}{\sqrt{d+1}}$. The appearance of $d-3$ in the conjecture accounts for $d=3$ being special in that $\Delta_{3}=0$. The values $d=1,2$ have $\Delta_{d}<0$, so $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$ is an imaginary quadratic field, while for $d\geq 4$, $\Delta_{d}$ specifies a (nontrivial) real quadratic field. The occurrence of a small unit $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$ (as compared to $\Delta_{d}$, which need not be a fundamental discriminant!) is certainly relevant. As noted by previously by the first author [32, p. 8], skipping repeated fields (e.g., $K_{8}=K_{4}$), the sequence $K_{4},K_{5},K_{6},K_{7},K_{9},K_{10},\ldots$ is precisely the ordering of all real quadratic fields by the size of the smallest nontrivial totally positive unit (which will necessarily be equal to $\varepsilon_{d}$).

### v3

@@ Abstract.
This paper concerns SIC-POVMs and their relationship to class field theory. SIC-POVMs are generalized quantum measurements (POVMs) described by $d^{2}$ equiangular complex lines through the origin in $\mathbb{C}^{d}$. Weyl–Heisenberg covariant SICs are those SIC-POVMs described by the orbit a single vector under a finite Weyl–Heisenberg group $\operatorname{WH}(d)$.
We relate known data on the structure and classification of Weyl–Heisenberg covariant SICs in low dimensions to arithmetic data attached to certain orders of real quadratic fields. For $4\leq d\leq 90$, we show the number of known geometric equivalence classes of Weyl–Heisenberg covariant SICs in dimension $d$ equals the cardinality of the ideal class monoid of (not necessarily invertible) ideal classes in the real quadratic order $\mathcal{O}_{\Delta}$ of discriminant $\Delta=(d+1)(d-3)$; we conjecture the equality extends to all $d\geq 4$. We prove that this conjecture implies the existence of more than one geometric equivalence class of Weyl–Heisenberg covariant SICs for every $d>22$. We conjecture that Galois multiplets of SICs may be put in one-to-one correspondence with the over-orders $\mathcal{O}^{\prime}$ of $\mathcal{O}_{\Delta}$ in such a way that the number of geometric classes in the multiplet equals the ring class number of $\mathcal{O}^{\prime}$. We test that conjecture against known data on exact SICs in low dimensions.
We refine the “class field hypothesis” of Appleby, Flammia, McConnell, and Yard [9] to predict the exact class field over $\mathbb{Q}(\sqrt{(d+1)(d-3)})$ generated by the ratios of vector entries for the equiangular lines defining a Weyl–Heisenberg SIC. The refined conjectures use a recently developed class field theory for orders of number fields [34]. The refined class fields assigned to over-orders $\mathcal{O}^{\prime}$ have a natural partial order under inclusion; the inclusions of these class fields fail to be strict in some cases. We characterize such cases and give a table of them for $d<500$.
@@
SIC-POVM, complex equiangular lines, equiangular tight frame (ETF), real quadratic field, non-maximal order, ideal class group, explicit class field theory
2020 Mathematics Subject Classification:
11R37 (primary), 11R29, 11R65, 81P15, 81P18, 81R05, 42C15
@@ 1. Introduction
SIC-POVMs are generalized quantum measurements (POVMs) described by $d^{2}$ equiangular complex lines through the origin in $\mathbb{C}^{d}$. Sets of $d^{2}$ equiangular complex lines in $\mathbb{C}^{d}$ are extremal objects in combinatorial design theory; they give equiangular tight frames [46, Ch. 12–14]. The name SIC-POVM (abbreviated ‘SIC’) was introduced in a 2004 paper of Renes, Blume-Kohout, Scott, and Caves [38], whose study was motivated by quantum state tomography, the problem of reconstructing a quantum state from a series of measurements performed on identical states. Scott showed that SICs are optimal measurement ensembles for quantum tomography [39]; see also [25]. SICs have found further applications to quantum information processing [13, 44, 45, 17], quantum foundations [24], compressed sensing for radar [29], and classical phase retrieval [21]. At present SICs have been constructed for finitely many dimensions $d$, and conjecturally they exist for all $d$.
The objects of this paper are to extend the range of number-theoretic conjectures about the existence and structure of SICs proposed by Appleby, Flammia, McConnell, and Yard [8, 9] and to test the refined conjectures against empirical data. We also draw consequences of the refined conjectures for counting and classifying equivalence classes of SICs. The conjectures of [9] relate Weyl–Heisenberg SICs in $\mathbb{C}^{d}$ (defined in Section 2) for $d\geq 4$ to class field theory over real quadratic fields. The ratios of entries of each of the $d^{2}$ vectors in the currently known SICs lie in a number field that is an abelian extension of $K=\mathbb{Q}(\sqrt{\Delta_{d}})$ with $\Delta_{d}=(d+1)(d-3)$. Class field theory classifies the abelian extensions of a given number field $K$ in terms of ray class groups, which encode the arithmetic of ideals of the ring $\mathcal{O}_{K}$ of algebraic integers of $K$.
The number fields associated to individual SICs in dimension $d$ may vary with the SIC, and the subclass of SICs to which the most precise conjecture of [9] applies consists of those giving rise to the minimal number field that occurs for a fixed $d$. Appleby, Flammia, McConnell, and Yard conjecture that this minimal field is a specific ray class field of $\mathbb{Q}(\sqrt{\Delta_{d}})$; see 1.1. The corresponding SICs comprise the minimal multiplet, as defined in Section 2. It is one of several Galois multiplets, which are finite sets of geometric equivalence classes of SICs making up a single orbit of a Galois group action.
The current paper formulates a refinement of these conjectures that accounts for all the remaining Galois multiplets of Weyl–Heisenberg SICs, those having larger associated number fields. For $d\geq 4$ the refined conjecture asserts that the multiplets can be indexed bijectively by the the orders $\mathcal{O}^{\prime}$ of the real quadratic number field $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$ intermediate between the order $\mathcal{O}=\mathbb{Z}[\varepsilon_{d}]$ with $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$ and the maximal order $\mathcal{O}_{K_{d}}$ of all algebraic integers in $K_{d}$. Here an order $\mathcal{O}$ of $K_{d}$ is a subring of $K_{d}$ which contains $1$ and is of finite index in $\mathcal{O}_{K}$.
In addition, the number field associated to the $\mathcal{O}^{\prime}$-multiplet is conjectured to be a certain ray class field associated to the order $\mathcal{O}^{\prime}$. The notion of ray class fields associated to orders is an extension of the usual class field theory (which is the special case of the maximal order), variants of which were developed (independently) by Campagna and Pengo [12] and the authors [34]. Inclusions of orders induce reverse inclusion of the corresponding ray class fields, so the minimal multiplet corresponds to the maximal order.
1.1. SICs
@@
A set $S$ of complex lines through the origin in $\mathbb{C}^{d}$, written $S=\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{n}\}\subset\mathbb{P}^{d-1}(\mathbb{C})$ with $\mathbf{v}_{j}$ unit vectors, is equiangular if all the pairwise Hermitian inner products $\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle$ for $i\neq j$ have the same absolute value $\left|\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle\right|=\alpha\geq 0$. In 1975, Delsarte, Goethals, and Seidel [18] proved an upper bound of $n\leq d^{2}$ on the size of any set of complex equiangular lines in $\mathbb{C}^{d}$. They showed that if the bound $n=d^{2}$ is attained, then necessarily $\alpha=\frac{1}{\sqrt{d+1}}$.
@@
A symmetric informationally complete positive operator-valued measure or SIC-POVM (henceforth SIC) is a set of normalized projections onto a (maximal) set of $d^{2}$ complex equiangular lines in $\mathbb{C}^{d}$. For $d^{2}$ equiangular lines $\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{d^{2}}\}$ represented by unit column vectors $\mathbf{v}_{j}\in\mathbb{C}^{d}$, the associated SIC is the set of normalized rank one Hermitian projection operators ${\mathbf{\Pi}}(S):=\left\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\right\}$ where $\Pi_{j}=\mathbf{v}_{j}\mathbf{v}_{j}^{\dagger}$. (Vectors $\mathbf{v}$ are column vectors, $\mathbf{v}^{\dagger}$ is the conjugate transpose of $\mathbf{v}$, and the Hermitian inner product is $\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle=\mathbf{v}_{i}^{\dagger}\mathbf{v}_{j}$, following the quantum mechanics convention established by Dirac [19, Sec. 1.6]. Here, $\mathbf{v}_{i}^{\dagger}\mathbf{v}_{i}=1$.) The SIC condition is equivalent to the following $d^{4}$ identities on the traces of products of the projections:
@@
We will use the term “SIC” interchangeably to mean either a “set of $d^{2}$ (scaled) Hermitian projection matrices ${\mathbf{\Pi}}(S)$ forming a SIC-POVM” or a “set of $d^{2}$ complex equiangular lines $S$ in $\mathbb{C}^{d}$.” When more specificity is needed, we term the latter a line-SIC.
@@
Most work on SICs follows Zauner’s 1999 PhD thesis [48, 49] on quantum designs. A few sporadic examples (such as the Hoggar lines in dimension $8$ [30, 31]) were known prior to Zauner’s work. Zauner conjectured, based on his numerical investigations, that SICs exist in every dimension $d$, and moreover, that a SIC in $\mathbb{C}^{d}$ could be found as the orbit of a single vector under the Weyl–Heisenberg group [48, 49]; the latter objects are called Weyl–Heisenberg covariant SICs (or just Weyl–Heisenberg SICs) and are defined and discussed in Section 2. A generating vector ${\mathbf{v}}\in\mathbb{C}^{d}$ of such an orbit is called a fiducial vector. Weyl–Heisenberg SICs are the main topic of this paper. (An extensive treatment of the history of SICs is given by Fuchs, Hoang, and Stacy [23].)
@@
Over the past 25 years, a large data set has been computed about SICs in particular dimensions—sometimes approximately and sometime exactly—resulting in extensive datasets [41, 40, 22]. SICs have been rigorously constructed in a finite set of dimensions, including Weyl–Heisenberg SICs in all dimensions $1\leq d\leq 53$ and the Hoggar lines for $d=8$. Beyond these exact constructions, numerical computations in dimensions including every $d\leq 90$ suggest there are only finitely many (equivalence classes of) Weyl–Heisenberg SICs in each dimension $d\neq 3$. It is known that $d=3$ has infinitely many equivalence classes of Weyl–Heisenberg SICs, and we will not treat $d=3$ here.
@@
The current numerical data is believed to account for all Weyl–Heisenberg SICs of dimension $d\leq 90$. There is a geometric action moving these SICs around by the normalizer of the Weyl–Heisenberg group inside the extended unitary group (which is a finite group whose order varies erratically with $d$, while the Weyl–Heisenberg group itself has order $d^{3}$ or $2d^{3}$, the latter occurring for even $d$.) In addition, since the (normalized) generating vectors of Weyl–Heisenberg SICs can be (empirically) given by algebraic numbers, there is a Galois action, which sometimes moves SICs to new SICs and sometimes moves them to non-SICs.
@@
    the finiteness of the set of Weyl–Heisenberg line-SICs in every fixed dimension $d\neq 3$.
The finiteness assertion has been a “folklore” conjecture since about 2004. It was observed empirically under heuristics up to dimension $d=7$ by Renes, Blume-Kohout, Scott, and Caves [38, Tab. I and Sec. V], up to dimension $d=50$ by Scott and Grassl (who state they are confident their list is complete for $d\leq 50$ in [41, Sec. 4]), and up to $d=90$ by Fuchs, Hoang, and Stacy [23, Sec. 6] using data of Scott [40]. It was explicitly formulated as a conjecture in the first author’s PhD thesis [32, Conj. V.12]. A closely related conjecture was formulated by Waldron [46, Sec. 14.27].
The finiteness assertion would imply fiducial vectors of Weyl–Heisenberg SICs can be chosen to have all entries algebraic numbers. Further study of these algebraic numbers in existing examples has led to surprising discoveries.
@@ 1.2. Connection of SICs to algebraic number theory and class fields
Recent work has connected SICs to explicit class field theory over real quadratic number fields. Appleby, Yadsan-Appleby, and Zauner [4] discovered in 2013 that in all known examples, the algebraic number field $\mathbb{E}$ generated by the entries of the projections ${\mathbf{\Pi}}(S)$ contains the quadratic field $K_{d}=\mathbb{Q}(\sqrt{(d+1)(d-3)})$, and Galois automorphisms often map SICs to new SICs. Moreover, they observed in all known examples (for $d\neq 3$) that $\mathbb{E}$ is an abelian extension of $K_{d}$.
In 2016, Appleby, Flammia, McConnell, and Yard [9, 8] made explicit predictions relating the number fields generated by Weyl–Heisenberg SICs to ray class fields. Based on extensive data given in [9, Prop. 1], they formulated the following conjecture [9, Conj. 2].
@@
    Every Weyl–Heisenberg line-SIC $S$ in dimension $d$ has SIC field $F^{\rm vec}\!\left(S\right)$ that is a finite Galois extension of $\mathbb{Q}$ containing $H_{d^{\prime}\infty_{1}\infty_{2}}$.
The technical definition of the SIC field $F^{\rm vec}\!\left(S\right)$ appearing in 1.1 is given in Section 3.1. Ray class fields are discussed in Section 6.1.
@@
They verified assertion (2) of the conjecture for all known Weyl–Heisenberg SICs in these dimensions. They also verified that the fields $F^{\rm vec}\!\left(S\right)$ are all abelian over $K_{d}$ in the known cases. (The list of SICs used is believed to be complete in these dimensions.)
@@
The origin of the discriminant $\Delta_{d}=(d+1)(d-3)$ appearing in the conjecture is not completely understood. The appearance of the factor $d+1$ dividing $\Delta_{d}$ seems related to the fact that the inner products of distinct unit vectors in a SIC are of absolute value $\frac{1}{\sqrt{d+1}}$. The appearance of $d-3$ accounts for $d=3$ being special, in that $\Delta_{3}=0$. The values $d=1,2$ have $\Delta_{d}<0$, so $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$ is an imaginary quadratic field, while for $d\geq 4$, $\Delta_{d}$ specifies a (nontrivial) real quadratic field. The occurrence of a small unit $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$ (as compared in size to $\Delta_{d}$, which need not be a fundamental discriminant) is certainly relevant. The first author previously noted that, on skipping repeated fields (e.g., $K_{8}=K_{4}$), the sequence $K_{4},K_{5},K_{6},K_{7},K_{9},K_{10},\ldots$ gives precisely the ordering of all real quadratic fields by increasing size of the smallest nontrivial totally positive unit (necessarily equal to $\varepsilon_{d}$); see [33, p. 13819].